Mass of an electron is X times to mass lf hydrogen atom.what is approximate value of X.
 

Grade:12

1 Answers

Arun
5 years ago
Dear Simran
Mass of an electron = 9.1 * 10^ – 31 kg
Mass of Hydrogen atom = 1.67 * 10^ –27 kg
 
hence 
X = 9.1 * 10^-31 / 1.67 * 10^-27
 
X = 5.44 * 10^-4
 
x = 0.000544